# High-probability Convergence Guarantees of Decentralized SGD

### v3

@@ 1.1 Literature Review
Centralized HP Convergence. [54] show optimal convergence rates of SGD for convex costs under light-tailed stochastic gradients (see assumption (A4) ahead for a formal definition of light-tailed noise), while [20] show the same for non-convex costs. [36, 37] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively. [21] and [5] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while [41] establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs. [39] generalize the previous works on non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs. [46] study HP convergence of SGD under sub-Weibull noise. Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [50, 56, 4] for various conditions used to study heavy-tailed noise. e.g., [61, 56, 40, 58, 23, 35, 3, 2, 4], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1). Crucially, convergence in the HP sense is achieved under the same conditions on the cost function as in the MSE sense, for both light-tailed and heavy-tailed noise.
@@
Decentralized HP Convergence. Compared to MSE guarantees, there is a significantly smaller body of work on HP convergence guarantees of decentralized algorithms. In particular, [44] study HP convergence of DSGD for general non-convex and PŁ costs under light-tailed noise, requiring uniformly bounded gradients, as well as asymptotically vanishing noise for PŁ costs (see the discussion after Theorem 2 for details), while showing optimal rates in both cases. [45, 76] study HP convergence of a decentralized mirror descent algorithm under light-tailed noise, for online noncooperative games and dynamic regrets, respectively, requiring bounded gradients and compact domains. [59, 77] study HP convergence of decentralized algorithms with clipping, for convex costs under heavy-tailed noise, while [78] also consider general non-convex costs. It is worth mentioning works like [7, 6, 47, 48, 8], where the authors study asymptotic large deviation guarantees for decentralized problems such as detection and inference. A common thread for all these works is the need for uniformly bounded gradients, or algorithmic modifications like gradient clipping, which ensure that the gradients stay bounded, as well as the fact that none of the existing HP studies achieve linear speed-up in the number of users, which is a known feature of decentralized algorithms in the MSE sense in, e.g., [33, 75, 1]. These observations raise the following important questions: (i) Can decentralized algorithms converge in the HP sense under the same assumptions on the cost as used for MSE results, where bounded gradients are not required? (ii) Is it possible to achieve a linear speed-up in the number of users in the HP sense, which is a well-known feature of decentralized algorithms? This situation is further contrasted by the centralized setting, where it is known that SGD-type algorithms require the same conditions on the cost function to guarantee both HP and MSE convergence.
@@ 3.1 Preliminaries
Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [74]. [^5] Our work can be readily extended to the setting of time-varying networks, see Appendix J for a detailed discussion. Let $\lambda\coloneqq\|W-J\|$ denote the network connectivity parameter, where $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$ is the ideal communication matrix. It can be shown that (A1) implies $\lambda\in[0,1)$, see, e.g., [22].
@@
Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [20, 75].
@@
    The noise at user $i\in[n]$ is $\sigma_{i}$-sub-Gaussian (light-tailed), i.e., for all $t\geq 1$ and $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
@@
The first condition in (A4) is standard in decentralized stochastic optimization [75, 79], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian). Light tails are necessary to achieve exponentially decaying tail bounds of the form in (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [54, 20, 36, 37, 21, 39, 5]. [^6] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [46] and references therein, the focus of this work is on sub-Gaussian noise.
The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
@@
    The average noise $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian, i.e., $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{120d\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1)$.
Lemma 2 provides some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, establishing the variance reduction benefit of decentralized optimization in the HP sense. This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we do in the following subsections. Note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and Appendix K.
@@ 3.2 Non-Convex Costs
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$
@@
$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
Theorem 1 establishes HP convergence of DSGD for general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few different perspectives.
@@
We can see from the bound in Theorem 1 that network connectivity only affects higher-order terms, decaying at rate $\mathcal{O}\Big(\frac{1}{T}\Big)$, which is consistent with MSE results, e.g., [33]. Moreover, the rate in Theorem 1 implies that the transient time of DSGD in the HP sense is of order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD implied by MSE rates, see, e.g., Table I in [1]. For details on the derivation of the transient time, see Appendix H.
@@
The authors in [44] study HP guarantees of DSGD over undirected time-varying networks, under a time-varying step-size schedule $\alpha_{t}=\frac{a}{\sqrt{t+1}}$, assuming uniformly bounded gradients and establish the rate $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$. As highlighted next, we provide several important improvements. First, we remove the uniformly bounded gradient assumption, replacing it with the more general heterogeneity condition (A5), which is facilitated by the use of the “offset trick”, see the discussion in Section 4 for details. Next, [44] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise. In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, see Appendix E for a detailed discussion. However, it would be impossible for [44] to obtain a linear speed-up even under a fixed step-size, as they show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity. On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used, see Appendix I for detailed discussion on the results needed to achieve linear speed-up. [^8] This is consistent with MSE results, e.g., [33, 75, 1], where linear speed-up in non-convex case is achieved only via a fixed step-size. Finally, while we consider static networks, our results can be extended to the time-varying setting in [44], see Appendix J for details.
@@ 3.3 Strongly Convex Costs
Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5), with $S_{k},D_{k}$ depending on the step-size and other problem parameters, see Appendix F for explicit forms of $S_{k},D_{k}$. We note that the requirement of same model initialization across users can be removed, at the cost of an additional, geometrically decaying term in the exponent on the RHS, however, we use same initialization, for ease of exposition. Prior to stating the main theorem, we provide an important technical result, which facilitates sharper bounds and ensures linear speed-up in the number of users.
@@
Lemma 5 provides a tight bound on the MGF of an “almost decreasing” process, with the optimality gap of strongly convex costs being an instance of such a process. Compared to bounds used in the centralized setting, e.g., [21, 5, 41], Lemma 5 is significantly sharper and incorporates higher-order terms, which are crucial for ensuring linear speed-up, see Section 4 and Appendix I for detailed discussions. We are now ready to state the main result for strongly convex costs.
@@
Let (A1)-(A4) and (A6) hold, let the step-size schedule be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If constants $a$ and $t_{0}$ are chosen as $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
$\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+d\sigma^{2}\kappa/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(T+t_{0})^{2}}\bigg),$
@@
Theorem 2 establishes HP convergence for smooth strongly convex costs and time-varying step-size. In Appendix F we provide the full bound, with additional higher-order terms. Next, we discuss multiple aspects of our results.
@@
We can again see that network connectivity only affects higher-order terms. Morevoer, the rate in Theorem 2 implies a transient time of DSGD in the HP sense, of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$. This is strictly sharper than the transient time of DSGD implied by MSE rates, of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see Table II in [1], further highlighting the tightness of our results and improved analysis.
@@
While heterogeneity only affects higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see Appendix F for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule. Similarly to the discussion after Theorem 1, the effect of noise on the leading term can be mitigated by employing a carefully selected fixed step-size. Due to time-varying step-sizes being sufficient for achieving linear speed-up and for ease of exposition, the results using a fixed step-size are omitted.
@@
where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size. Both conditions are difficult to satisfy, as $(i)$ is a path-wise relaxation of the uniformly bounded gradient condition (which is not satisfied for strongly convex costs), while $(ii)$ implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that it vanishes at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$. [^9] Markov’s inequality implies that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\Big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\Big)\geq 1-\delta$, for any $\delta\in(0,1)$. Similarly, Jensen’s inequality implies $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$. Crucially, [44] are unable to establish linear speed-up, even though they impose much stronger assumptions on the noise and cost. On the other hand, we show optimal convergence rate with linear speed-up using standard assumptions, with sharper transient times compared to MSE results. These improvements stem not only from a tighter analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity condition (A5) and ensure linear speed-up is achieved, see Appendix I for a further discussion.
@@
In Appendix G we provide numerical experiments which validate the main findings of the paper, notably showing that (i) DSGD converges in HP under the same conditions on the cost as in the MSE sense and (ii) DSGD achieves linear speed-up in the HP sense. The reader is kindly referred to the Appendix for details.

### v5

@@ 1.1 Literature Review
Centralized HP Convergence. [58] show optimal convergence rates of SGD for convex costs under light-tailed stochastic gradients (see assumption (A4) ahead for a formal definition of light-tailed noise), while [20] show the same for non-convex costs. [39, 40] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively. [23] and [5] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while [44] establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs. [42] generalize the previous works on non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs. [49] study HP convergence of SGD under sub-Weibull noise. Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [53, 60, 4] for various conditions used to study heavy-tailed noise. e.g., [21, 68, 60, 43, 63, 25, 37, 3, 2, 4], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1). Crucially, convergence in the HP sense is achieved under the same conditions on the cost function as in the MSE sense, for both light-tailed and heavy-tailed noise.
@@
Decentralized HP Convergence. Compared to MSE guarantees, there is a significantly smaller body of work on HP convergence guarantees of decentralized algorithms. In particular, [47] study HP convergence of DSGD for general non-convex and PŁ costs under light-tailed noise, requiring uniformly bounded gradients, as well as asymptotically vanishing noise for PŁ costs (see the discussion after Theorem 2 for details), while showing optimal rates in both cases. [48, 83] study HP convergence of a decentralized mirror descent algorithm under light-tailed noise, for online noncooperative games and dynamic regrets, respectively, requiring bounded gradients and compact domains. [64, 84] study HP convergence of decentralized algorithms with clipping, for convex costs under heavy-tailed noise, while [85] also consider general non-convex costs. Finally, [22] study clipped SGD under heavy-tailed noise, in the distributed client-server (i.e., fully connected network) setting. It is worth mentioning works like [7, 6, 50, 51, 8], where the authors study asymptotic large deviation guarantees for decentralized problems such as detection and inference. A common thread for all these works is the need for uniformly bounded gradients, or algorithmic modifications like gradient clipping, which ensure that the gradients stay bounded, as well as the fact that none of the existing HP studies achieve linear speed-up in the number of users, which is a known feature of decentralized algorithms in the MSE sense in, e.g., [35, 82, 1]. These observations raise the following important questions: (i) Can decentralized algorithms converge in the HP sense under the same assumptions on the cost as used for MSE results, where bounded gradients are not required? (ii) Is it possible to achieve a linear speed-up in the number of users in the HP sense, which is a well-known feature of decentralized algorithms? This situation is further contrasted by the centralized setting, where it is known that SGD-type algorithms require the same conditions on the cost function to guarantee both HP and MSE convergence.
@@ 3.1 Preliminaries
Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [81]. [^5] Our work can be readily extended to the setting of time-varying networks, see Appendix K for a detailed discussion. Let $\lambda\coloneqq\|W-J\|$ denote the network connectivity parameter, where $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$ is the ideal communication matrix. It can be shown that (A1) implies $\lambda\in[0,1)$, see, e.g., [24].
@@
Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [20, 82]. While some works, e.g., [38], assume gradients of loss $\ell$ are $L$-Lipschitz continuous, we note that while similar, such assumption and (A3) aim to exploit different structural properties. However, it is known that under certain regularity conditions, $L$-Lipschitz gradients of $\ell$ imply $L$-Lipschitz gradients of $f_{i}$, making such condition slightly stronger than (A3).
@@
    The gradient noise $z^{t}_{i}$ at user $i\in[n]$ and time $t\geq 1$ is $\sigma_{i}$-sub-Gaussian, i.e., we have
@@
The first condition in (A4) is standard in decentralized stochastic optimization [82, 86], while the second and third require noise to be unbiased and sub-Gaussian (i.e., light-tailed). Light tails are necessary to achieve exponentially decaying tail bounds of the form in (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [58, 20, 39, 40, 23, 42, 5]. [^6] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [49] and references therein, the focus of this work is on sub-Gaussian noise.
The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, [^7] With at least one of $A,B$ being strictly positive. we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
@@
    The average noise $\overline{z}^{t}$ satisfies $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{15\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp\big(1+\log(2d)\big)$.
Lemma 2 provides some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma}{\sqrt{n}}\Big)$-sub-Gaussian, establishing the variance reduction benefit of decentralized optimization in the HP sense. This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we do in the following subsections. We note that the third property introduces a mild logarithmic dependence on the problem dimension $d$, with further discussion provided in Section 4 and Appendix L.
@@ 3.2 Non-Convex Costs
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{15LT}}\right\}$, where $C>0$ is a problem related constant satisfying
$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{3\sigma\sqrt{5L}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$
@@
$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{L}\big(\Delta_{f}+\log(\nicefrac{{2d}}{{\delta}})\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}(1-\lambda)^{2}T}\Bigg).$
Theorem 1 establishes HP convergence of DSGD to a stationary point (or a set of stationary points), for general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few different perspectives.
@@
We can see from the bound in Theorem 1 that network connectivity only affects higher-order terms, decaying at rate $\mathcal{O}\Big(\frac{1}{T}\Big)$, which is consistent with MSE results, e.g., [35]. Moreover, the rate in Theorem 1 implies that the transient time of DSGD in the HP sense is of order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD implied by MSE rates, see, e.g., Table I in [1]. For details on the derivation of the transient time, see Appendix I.
@@
The authors in [47] study HP guarantees of DSGD over undirected time-varying networks, under a time-varying step-size schedule $\alpha_{t}=\frac{a}{\sqrt{t+1}}$, assuming uniformly bounded gradients and establish the rate $\mathcal{O}\Big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\Big)$. As highlighted next, we provide several important improvements. First, we remove the uniformly bounded gradient assumption, replacing it with the more general heterogeneity condition (A5), which is facilitated by the use of the “offset trick”, see the discussion in Section 4 for details. Next, [47] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise. In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, see Appendix F for a detailed discussion. However, it would be impossible for [47] to obtain a linear speed-up even under a fixed step-size, as they show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity. On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used, see Appendix J for detailed discussion on the results needed to achieve linear speed-up. [^9] This is consistent with MSE results, e.g., [35, 82, 1], where linear speed-up in non-convex case is achieved only via a fixed step-size. Finally, while we consider static networks, our results can be extended to the time-varying setting in [47], see Appendix K for details.
@@ 3.3 Strongly Convex Costs
Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5), with $S_{k},D_{k}$ depending on the step-size and other problem parameters, see Appendix G for explicit forms of $S_{k},D_{k}$. We note that the requirement of same model initialization across users can be removed, at the cost of an additional, geometrically decaying term in the exponent on the RHS, however, we use same initialization, for ease of exposition. Prior to stating the main theorem, we provide an important technical result, which facilitates sharper bounds and ensures linear speed-up in the number of users.
@@
Lemma 5 provides a tight bound on the MGF of an “almost decreasing” process, with the optimality gap of strongly convex costs being an instance of such a process. Compared to bounds used in the centralized setting, e.g., [23, 5, 44], Lemma 5 is significantly sharper and incorporates higher-order terms, which are crucial for ensuring linear speed-up, see Section 4 and Appendix J for detailed discussions. We are now ready to state the main result for strongly convex costs.
@@
Let (A1)-(A4) and (A6) hold, let the step-size schedule be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If constants $a$ and $t_{0}$ are chosen as $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{1960\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
$\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}\kappa\big(1+\log(2d)\big)/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\nicefrac{{\|\nabla\mathbf{f}^{\star}\|^{2}(1+\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(T+t_{0})^{2}}\bigg),$
@@
Theorem 2 establishes HP convergence of DSGD to the global minumum, for smooth strongly convex costs and time-varying step-size. In Appendix G we provide the full bound, with additional higher-order terms. Next, we discuss multiple aspects of our results.
@@
We can again see that network connectivity only affects higher-order terms. Moreover, the rate in Theorem 2 implies a transient time of DSGD in the HP sense, of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$. This is strictly sharper than the transient time of DSGD implied by MSE rates, of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see Table II in [1], further highlighting the tightness of our results and improved analysis.
@@
While heterogeneity only affects higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see Appendix G for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule. Similarly to the discussion after Theorem 1, the effect of noise on the leading term can be mitigated by employing a carefully selected fixed step-size. Due to time-varying step-sizes being sufficient for achieving linear speed-up and for ease of exposition, the results using a fixed step-size are omitted.
Extending to slower decaying step-size and Polyak-Ruppert average.
While Theorem 2 provides the optimal convergence rate for strongly convex costs, it is also of interest to consider step-sizes that decay at a slightly slower rate, e.g., $\alpha_{t}\propto(t+t_{0})^{-\eta}$, for $\eta\in[1/2,1)$. The techniques used in our proof can readily handle such step-sizes, but would inevitably lead to sub-optimal rates of order $\mathcal{O}\Big(\frac{1}{nt^{\eta}}\Big)$, while noting that the step-size schedule $\alpha_{t}\propto(t+t_{0})^{-1}$ used in Theorem 2 is known to be optimal for strongly convex costs, e.g., [66, 44]. Additionally, for (strongly) convex costs, it is often of interest to provide convergence guarantees in terms of the Polyak-Ruppert average model [67, 61]. Such guarantees readily follow, e.g., from Theorem 1 and using (A6). Indeed, defining the Polyak-Ruppert average as $\widetilde{x}^{t}=\frac{1}{nT}\sum_{i\in[n],t\in[T]}x^{t}_{i}$, using the results of Theorem 1 with step size $\alpha\propto\frac{1}{\sqrt{nT}}$ and (A6), we get
$2\mu\big(f(\widetilde{x}^{t})-f^{\star})\leq\frac{1}{nT}\sum_{i\in[n],t\in[T]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\bigg(\frac{1}{\sqrt{nT}}\bigg).$
However, we note that the guarantees of Theorem 2 are stronger in multiple facets: (i) the step-size and rate in Theorem 2 are optimal for strongly convex costs; (ii) the last iterate is practically more viable than the Polyak-Ruppert average and does not require the users to maintain an extra variable (i.e., the running average); (iii) the analysis for the last iterate is significantly more involved, as demonstrated in the appendix, whereas the result for Polyak-Ruppert average follow directly from Theorem 1 and the properties of strong convexity.
@@
where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size. Both conditions are difficult to satisfy, as $(i)$ is a path-wise relaxation of the uniformly bounded gradient condition (which is not satisfied for strongly convex costs), while $(ii)$ implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that it vanishes at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$. [^10] Markov’s inequality implies that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\Big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\Big)\geq 1-\delta$, for any $\delta\in(0,1)$. Similarly, Jensen’s inequality implies $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$. Crucially, [47] are unable to establish linear speed-up, even though they impose much stronger assumptions on the noise and cost. On the other hand, we show optimal convergence rate with linear speed-up using standard assumptions, with sharper transient times compared to MSE results. These improvements stem not only from a tighter analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity condition (A5) and ensure linear speed-up is achieved, see Appendix J for a further discussion.
@@
In Appendix H we provide numerical experiments which validate the main findings of the paper, notably showing that (i) DSGD converges in HP under the same conditions on the cost as in the MSE sense and (ii) DSGD achieves linear speed-up in the HP sense. The reader is kindly referred to the Appendix for details.